### v2

# Where Giants Dwell: Probing the Environments of Early Massive Quiescent Galaxies

11institutetext: INAF - Astronomical Observatory of Trieste, via G.B. Tiepolo 11, I-34143 Trieste, Italy 22institutetext: IFPU - Institute for Fundamental Physics of the Universe, via Beirut 2, 34151, Trieste, Italy 33institutetext: Tianjin Normal University, Binshuixidao 393, 300387, Tianjin, People’s Republic of China 44institutetext: Institute for Physics, Laboratory for Galaxy Evolution, EPFL, Observatoire de Sauverny, Chemin Pegasi 51, 1290 Versoix, Switzerland
Where Giants Dwell: Probing the Environments of Early Massive Quiescent Galaxies
Gabriella De Lucia, e-mail:gabriella.delucia@inaf.it    Lizhi Xie    Michaela Hirschmann    Fabio Fontanot
(Received ???, 2025; accepted ???, 2025)
We investigate the environments of massive quiescent galaxies at $3\lesssim z\lesssim 5$ using the GAlaxy Evolution and Assembly (GAEA) theoretical model. We select galaxies with $10^{10.8}\,{\rm M}_{\sun}$ and specific star formation rate below $0.3\times t^{-1}_{\rm Hubble}$, yielding in a sample of about $5,000$ galaxies within a simulated volume of $\sim 685$ Mpc. These galaxies have formation times that cover well the range inferred from recent observational data, including a few rare objects with very short formation time-scales and early formation epochs. Model high-z quiescent galaxies are $\alpha$-enhanced and exhibit a wide range of stellar metallicity, in broad agreement with current observational estimates. Massive high-z quiescent galaxies in our model occupy a wide range of environments, from void-like regions to dense knots at the intersections of filaments. Quiescent galaxies in under-dense regions typically reside in halos that collapsed early and grew rapidly at high redshift, though this trend becomes difficult to identify observationally due to large intrinsic scatter in star formation histories. The descendants of high-z massive quiescent galaxies display a broad distribution in mass and environment by $z=0$, reflecting the stochastic nature of mergers. About one-third of these systems remain permanently quenched, while most rejuvenation events are merger-driven and more common in overdense regions. Our results highlight the diversity of early quiescent galaxies and caution against assuming that all such systems trace the progenitors of present day most massive clusters.
Key Words.:
galaxies: formation – galaxies: evolution – galaxies: high-redshift – galaxies: star formation – galaxies: stellar content – galaxies: abundances
1 Introduction
The assembly and quenching of massive galaxies are central questions in modern astrophysics. In the local Universe, where star formation can be measured accurately for large samples of galaxies using a variety of indicators, galaxies are found to follow a ‘bimodal’ distribution (e.g. Blanton et al., 2003; Kauffmann et al., 2003; Baldry et al., 2004). This bimodality exhibits a well defined red sequence dominated by low star formation rates, and a blue cloud populated by galaxies that are still actively forming stars at non negligible rates. The wealth of observational information available in the local Universe has allowed us to characterize how the distribution in the sSFR/color versus stellar mass plane depends on the environment, defined either in terms of local galaxy density or through a more ‘global’ environmental proxy such as halo mass or large-scale over-density. The fraction of passive galaxies is found to be larger in over-dense regions (e.g. Kauffmann et al., 2004; Balogh et al., 2004; Baldry et al., 2006), indicating that environmental processes had a significant impact on at least a fraction of the quenched galaxy population observed today.
At earlier cosmic epochs, accurate measurements of the star formation are more difficult to obtain for statistical samples of galaxies, and a less accurate environmental characterization is typically possible due for example to only photometric redshift information available or sparse spectroscopic sampling rates. Yet, available data lead to general trends that are similar to those observed in the local Universe, with an overall decrease of the fraction of quiescent galaxies at earlier cosmic epochs (Brammer et al., 2009; Whitaker et al., 2011; Muzzin et al., 2013; Weaver et al., 2023). The measured quiescent fractions remain very large among massive galaxies in clusters up to $z\sim 1$ (e.g. Mei et al., 2009; van der Burg et al., 2020), and quiescent galaxies are still found in the cores of massive systems up to $z\sim 2$ albeit with a relatively large cluster-to-cluster variation (Andreon et al., 2014; Strazzullo et al., 2019; Willis et al., 2020).
The recent advent of the James Webb Space Telescope (JWST) has significantly advanced our ability to identify and characterize quiescent galaxies at very early cosmic epochs. Thanks to its unprecedented sensitivity and resolution in the near- and mid-infrared, deep surveys such as the Cosmic Evolution Early Release Science programme (CEERS - Finkelstein et al. 2023), and the JWST Advanced Deep Extragalactic Survey (JADES - Eisenstein et al. 2023) have revealed a population of massive quiescent galaxies that significantly outnumber expectations from theoretical models (Carnall et al., 2023; Valentino et al., 2023; Baker et al., 2025a, b, and references therein). While the number of spectroscopically confirmed massive quiescent galaxies at early epochs increases steadily, their environments remains poorly constrained from the observational point of view. A few studies report evidence of local galaxy overdensities around massive quiescent galaxies at $z\gtrsim 3$ (Kubo et al., 2021; McConachie et al., 2022; Ito et al., 2023; Jin et al., 2024), suggesting that at least a fraction of these can be found in overdense regions that potentially trace the earliest cluster progenitors. Given the challenges involved in a robust determination of the environment at these redshifts, it remains unclear if these few existing case studies are representative of the global population.
From a theoretical perspective, a strong correlation with environment would align with expectations from a biased hierarchical formation scenario in which early quenching occurs preferentially in the densest regions of the cosmic web. However, it remains an open question whether environmental processes are causally linked to quenching at these early times, or whether environment merely correlates with early halo collapse while internal mechanisms dominate the suppression of star formation. The latter scenario is more likely to apply to the most massive galaxies, that are expected to experience only internal physical processes for most of their lifetime (De Lucia et al., 2012).
A few recent studies have examined the environmental properties of quenched high-z massive galaxies within theoretical models of galaxy formation, and the results are somewhat contradictory. Kimmig et al. (2025) use the Magneticum Pathfinder111www.magneticum.org simulation and find that the 36 quiescent galaxies with stellar mass larger than $3\times 10^{10}\,{\rm M}_{\odot}$ identified at $z\sim 3.4$ reside in underdense regions relative to galaxies of similar stellar mass. Kimmig et al. (2025) argue that, in the framework of the physical model for stellar and AGN feedback implemented in Magneticum, this is a consequence of more efficient gas ejection and lack of new gas replenishment in underdense regions. In contrast, Kurinchi-Vendhan et al. (2024, see also ), using the IllustrisTNG [^2] https://www.tng-project.org/ simulations, find that the $\sim 70$ massive [^3] They select only galaxies with stellar mass larger than $10^{10.6}\,{\rm M}_{\odot}$ quiescent galaxies identified at $z\sim 3.7$ in the TNG100 and TNG300 volumes are preferentially located in more massive halos and denser regions. These regions provide larger initial reservoirs of infalling gas, which promotes black hole mass growth and leads to quenching via AGN feedback [^4] It is the radiatively inefficient regime that leads, in the framework of this model, to an efficient momentum injection by AGN feedback. TNG assumes a switch from the radiatively efficient to the inefficient mode that depends on the BH mass. Therefore, the faster black hole growth, promoted in over-dense regions, leads to earlier gas accretion in the inefficient regime.. Similar trends are reported in Szpila et al. (2025) for the SIMBA-C simulation. These studies suggest that investigating the environmental properties of massive quiescent galaxies at early cosmic epochs can offer valuable insights into their formation pathways, and into the physical processes shaping galaxy evolution in the early Universe.
In this study, we carry out a throughout study of the local and large-scale environment of massive quiescent galaxies at $3<z<5$ taking advantage of our GAlaxy Evolution and Assembly (GAEA) model. This is a state-of-the-art theoretical model that includes an updated treatment for both stellar and AGN feedback and, as we have shown in previous work, reproduces the observed quiescent fractions as well as number densities of massive quiescent galaxies out to $z\sim 3-4$. It therefore represents an ideal tool for the detailed study presented below. The layout of the paper is as follows: in Section 2, we introduce the GAEA model and the simulation used in this study. In Sections 3 and 4, we characterize the environment of high-z quiescent model galaxies and in Section 5 we analyze to what extent the physical properties of high-z quiescent galaxies depend on their environment. In Section 6, we characterize the future evolution of galaxies in our model sample and its dependence on the environment. Finally, we summarize our results and give our conclusions in Section 7.
2 The simulation and the galaxy formation model
GAEA [^5] Details about the model, and access to a selection of data products, can be found at: https://sites.google.com/inaf.it/gaea/ is a state-of-the-art theoretical model of galaxy formation that couples prescriptions for the evolution of different baryonic components with substructure based merger trees extracted from high-resolution cosmological N-body simulations. The model builds on that presented in the original work by De Lucia and Blaizot (2007), but it has been updated significantly over the past few years. Topical developments have been:
  1. (i)
    a detailed treatment of the non-instantaneous recycling of gas, energy, and metals, which enables the tracing of individual metal abundances (De Lucia et al., 2014);
  2. (ii)
    an updated parametrization of stellar feedback partially based on results from hydro-dynamical simulations, tuned to reproduce the observed galaxy stellar mass function up to $z\sim 3$ (Hirschmann et al., 2016);
  3. (iii)
    a treatment for partitioning the cold gaseous phase of model galaxies in its atomic and molecular components, tuned to reproduce the measured HI and H2 galaxy mass functions in the local Universe (Xie et al., 2017);
  4. (iv)
    a careful tracing of the angular momentum exchanges between different components, and a treatment for the non instantaneous stripping of the cold and hot gaseous components associated with satellite galaxies (Xie et al., 2020);
  5. (v)
    an improved model for cold gas accretion on super massive black holes (BHs) and AGN driven outflows that is tuned to reproduce the measured evolution of the AGN luminosity function up to $z\sim 4$ (Fontanot et al., 2020).
In previous work, we have shown that our model is able to reproduce a broad range of observational measurements including the observed multiphase gas content of central and satellite galaxies (Xie et al., 2020), the observed secondary dependence of the local galaxy mass-gas metallicity relation (De Lucia et al., 2020), and its evolution to higher redshift (Hirschmann et al., 2016; Fontanot et al., 2021). Notably for this work, the latest GAEA rendition (De Lucia et al., 2024; Xie et al., 2024) reproduces nicely the observed galaxy quenched fractions and the estimated number densities of massive quiescent galaxies (when accounting for cosmic variance) up to $z\sim 3-4$. As discussed in De Lucia et al. (2024, see also ), GAEA also performs significantly better than other recently published theoretical models in reproducing the estimated quenched fractions in galaxy clusters at $z\sim 1$.
The results presented in this paper are based on the Millennium Simulation (Springel et al., 2005). This simulation follows 2,1603 dark matter particles in a box of 500 ${\rm Mpc}\,{\rm h}^{-1}$ on a side, and assumes cosmological parameters consistent with WMAP1 ($\Omega_{\Lambda}=0.75$, $\Omega_{m}=0.25$, $\Omega_{b}=0.045$, $n=1$, $\sigma_{8}=0.9$, and $H_{0}=73\,{\rm km\,s^{-1}\,Mpc^{-1}}$). More recent estimations provide slightly different values for a few of these parameters, and in particular: a larger value for $\Omega_{m}$ and a lower value for $\sigma_{8}$. However, we do not expect these differences to affect significantly our model predictions and, in fact, our recent work based on the P-Millennium Simulation (Fontanot et al., 2025) has shown that results from the latest rendition of our model converge very well when run on different simulations from the Millennium suite (see also Wang et al., 2008; Guo et al., 2013).
For the analysis presented below, we have selected massive quiescent galaxies at $z\gtrsim 3.1$ and $z\lesssim 5.3$. Specifically, we have considered all model galaxies in the redshift range of interest with stellar mass larger than $10^{10.8}\,{\rm M}_{\sun}$ [^6] The stellar mass cut adopted is rather arbitrary and a compromise between the large estimated stellar mass quoted in recent observational studies (e.g. Carnall et al., 2024) and the goal to have a large enough statistical sample of model galaxies., after accounting for an observational uncertainty of 0.25 dex, and with specific star formation rate (sSFR) lower than $0.3\times t^{-1}_{\rm Hubble}$, where $t_{\rm Hubble}$ is the age of the Universe at the corresponding redshift. We note that the assumed uncertainty is rather conservative at these redshift, and that a different selection would not change qualitatively the results presented in this work. Whenever appropriate, we will discuss in the following how results depend on this convolution. We find a total of 4788 galaxies in the entire Millennium Simulation box satisfying these criteria. We do not prevent the descendant/progenitor of the same galaxies to be counted at different snapshots. About 87 per cent of the selected galaxies are centrals and about 7 per cent are orphans, i.e. galaxies whose parent dark matter substructure has been tidally stripped below the resolution limit of the simulation. In our model, these galaxies survive for a residual merger time before merging (this is justified by the fact that baryons are much more concentrated than dark matter). Since the number density of massive quiescent galaxies decreases rapidly with increasing redshift, the sample is dominated by galaxies at the lowest end of the redshift range considered: 2195 galaxies are found at $z\sim 3.1$, 1287 at $z\sim 3.3$ and 668 at $z\sim 3.6$. There are only 12 massive quiescent galaxies at $z\sim 5.3$ - the snapshot at the highest redshift considered.
As a term of comparison for part of the analysis presented below, we also consider a corresponding star forming sample, i.e. galaxies that are above the same stellar mass cut but have sSFR larger than the threshold assumed for quiescent galaxies. Our star forming sample is composed of 23,656 galaxies; $\sim 88$ per cent of these are centrals and only $\sim 4$ per cent are orphans.
3 Are the first quiescent galaxies sitting in the most massive halos?
Figure 1: Top panel: relation between stellar mass and parent halo virial mass for all galaxies selected at $z\gtrsim 3.1$ and $z\lesssim 5.3$ (trends are similar when considering galaxies in smaller redshift ranges). Gray and cyan contours show the regions enclosing (from thicker to thinner lines) 30, 60, and 90 per cent of the quiescent and star forming galaxies, respectively. Bottom panel: same as above but including an estimate of the observational uncertainty ($0.25$ dex) in the galaxy stellar mass.
The first question we address relates to which halos host the most massive quiescent galaxies at $z\gtrsim 3$. The naive expectation (but see discussion in Section 1) is that these galaxies can be considered as the signposts of significant overdensities in the early Universe, that might eventually collapse in the most massive systems at lower redshifts. However, the stellar-to-halo mass relation tends to be rather flat for halo masses larger than $\sim 10^{13}\,{\rm M}_{\odot}$ (Girelli et al., 2020, and references therein), suggesting that a broader distribution of halo masses should be expected. Fig. 1 shows the correlation between galaxy stellar mass and parent halo mass for all model galaxies in our samples. Gray and cyan contours show the regions enclosing (from thicker to thinner lines) 30, 60, and 90 per cent of the quiescent and star forming galaxies, respectively. To compute the distributions shown, we have combined galaxies in the entire redshift range considered ($3\lesssim z\lesssim 5.3$), but we have verified that the trends are the same when considering narrower redshift bins. The figure shows that massive quiescent galaxies reside in halos that cover a broad range of masses, from Milky-Way like systems to the most massive systems that can be identified at these high redshifts. In this sense, they are not ‘special’ when compared to star forming galaxies of similar mass, that cover a similar range of parent halo masses. However, the distributions obtained for star forming galaxies tend to be shifted to lower halo masses with respect to those corresponding to quiescent galaxies, and the most massive halos tend to host primarily quiescent galaxies. This is in agreement with findings from Kurinchi-Vendhan et al. (2024) based on the IllustrisTNG simulation, and can be understood as a consequence of an accelerated growth of galaxies sitting at the center of very massive halos, and suppression of their star formation activity either because of mergers or because of AGN feedback. In our previous work, we have shown that the dominant quenching channel for massive galaxies in GAEA is, in fact, AGN feedback (Xie et al., 2024). Similar conclusions have been reached by independent studies based on both hydro-dynamical simulations and semi-analytic models (Kurinchi-Vendhan et al., 2024; Szpila et al., 2025; Lagos et al., 2025), with the exception of the study by Kimmig et al. (2025) mentioned above and based on the Magneticum simulation.
Fig. 1 shows that there is a weak correlation between the intrinsic galaxy stellar mass and parent halo mass (top panel), both for quiescent and star forming massive galaxies. However, such a correlation is completely washed out when convolving the intrinsic galaxy mass with a (rather optimistic) estimate of the observational uncertainties (bottom panel). Being rather flat as a function of virial mass, the distributions shown in the bottom panel of Fig. 1 do not vary significantly when further including an estimate for the observational uncertainty on the parent halo mass.
Figure 2: Halo mass distribution for quiescent model galaxies (solid lines) and the corresponding halo mass function (dashed lines). Different colors correspond to different redshifts (as indicated in the legend in the bottom panel), ranging from $\sim 3.1$ (blue) to $\sim 5.3$ (brown). The top and bottom panels are for all quiescent galaxies and for quiescent satellites only, respectively.
Figure 3: Spatial distribution of galaxies selected at $z\sim 4.5$ and with over-densities over a 3 Mpc scale in the lowest 15th (top panels) and highest 85th (bottom panels) percentiles of the distribution. Galaxies in our sample of massive quiescent galaxies are at the center of each panel and are marked by a red cross, while the other symbols mark the position of all galaxies more massive than $10^{9}\,{\rm M}_{\sun}$ in a projected region of 100x100 Mpc comoving with a dept of $20$ Mpc comoving. The o33 value corresponding to each field is indicated in the top left corner of each box. The color of the symbols scales with the galaxy star formation rate as indicated by the color bar on the right, while symbol size scales with galaxy stellar mass. We note that we have plotted the most massive galaxies at the end to emphasize their position, so they can hide other galaxies that lie close in projection.
Fig. 2 shows, for all cosmic epochs considered, how the distributions of halo masses corresponding to the massive quiescent galaxies in our sample (solid lines) compare to the halo mass functions at the same cosmic epoch (dashed lines). Different colors correspond to different redshifts, from $\sim 3.1$ (blue) to $\sim 5.3$ (brown). The top and bottom panels are for all quiescent galaxies and for quiescent satellite galaxies in our sample, respectively. The figure confirms that, while the most massive halos at each cosmic epoch tend to host central quiescent galaxies, many central galaxies in halos more massive than $\sim 10^{13}\,{\rm M}_{\sun}$ are forming stars at non negligible rates and therefore are not counted in the solid histograms in the top panel of Fig. 2. The few satellite galaxies ($\sim 620$ out of 4788) in our sample do not exclusively reside in the most massive halos but are distributed in the entire halo mass range covered by our sample. Combining the two panels, the figure also shows that the fraction of quiescent galaxies increases with halo mass, as observed at least in the local Universe (for a quantitative comparison between our model predictions and observational data, see De Lucia et al., 2019), and that this is not simply due to an increase of the satellite fraction.
4 What is the environment of high-z quiescent galaxies?
A measurement of parent halo mass is generally not easily accessible, in particular for the massive quiescent systems that are being observed at $z>3$. In this section, we introduce different estimates that might be more easily obtained from observational data. In particular, for each galaxy in our sample, we compute:
  • (i)
    a ‘local density’ estimate corresponding to the density of galaxies inside the projected circular region enclosing the 3 closest neighbors. We refer to this density as n33 in the following;
  • (ii)
    an ‘over-density’ estimate corresponding to the mean over-density of galaxies measured within a projected circular region of radius equal to 3 Mpc. This is normalized to the average value obtained at each redshift, by randomly drawing 1000 spheres with radius equal to 3 Mpc in the entire volume of the simulation. We refer to this over-density as o33 in the following.
In both cases, we have considered projected physical distances and only galaxies in slices of depth $\sim 20\penalty 10000\$Mpc with stellar mass larger than $10^{9}\,{\rm M}_{\sun}$. We have verified that the results discussed below do not vary significantly when increasing the number of neighbors (up to 5 or 7) or the scale considered (we have considered overdensities on scales of 5 and 8 Mpc).
For illustrative purposes, we show in Fig. 3 the projected spatial distribution of eight galaxies, all at $z\sim 4.5$, from our sample of quiescent galaxies. The galaxies have been selected to emphasize the variety of environments in which model high-z quiescent galaxies live. The top panels correspond to four galaxies in the lowest 15th percentile of the distribution of the over-densities over 3 Mpc, while the bottom panels correspond to galaxies in the highest 85th percentiles of the same distribution. Selected galaxies are marked by a red cross and are shown at the center of the boxes, that correspond to a depth of $20$ Mpc along the line of sight. We show the projected distribution of all galaxies more massive than $10^{9}\,{\rm M}_{\sun}$ at the same redshift; symbol size scales with galaxy stellar mass while color scales with the galaxy star formation rate, as indicated by the color bars on the right of the figure. We note that we have plotted the most massive galaxies at the end, so they might hide smaller galaxies at close projected distances - this happens for example in the third bottom panel where the massive quiescent galaxy at the center of the box is behind another massive star forming galaxy (see Figs. 11 and 12 in Appendix A for a similar plot, but showing only the distribution of the galaxies with low and high star formation rates in the same regions). The mean overdensity, which is indicated in the top left corner of each box, is increasing from left to right in each row.
Figure 4: Evolution of the parent halo mass, galaxy stellar mass, star formation rate, and black hole mass for the massive quiescent galaxies at $z\sim 4.5$ considered in Fig. 3. Green and magenta lines refer to the top and bottom panel respectively, with thickness of the lines and color shade increasing with over-density values as indicated in the legend included in the left panel. Different line styles are used to show different galaxies for clearer visualization. $\rm{M}_{200}$ and $\rm M_{BH}$} are shown for the main progenitor at each previous cosmic epoch, while for the galaxy stellar mass and SFR we show the sum of all progenitors at the corresponding epoch.
Figure 5: Correlation between two different estimates of the galaxy environment and the parent halo mass (left column), the intrinsic galaxy stellar mass (middle column), and the galaxy stellar including an estimate for its uncertainty (right column). The top panels are for n33, i.e. the local density (in physical Mpc-2) estimated considering the three closest neighbors for each galaxy in our quiescent (black) and star forming (cyan) samples. The bottom panels are for o3, i.e. the over-density estimated considering a scale of 3 Mpc. The numbers in each panel correspond to the Spearman’s correlation coefficient and relative significance (a small value indicates a significant correlation) for quiescent galaxies (top left) and for star forming galaxies (bottom right).
Fig. 3 shows that there is a wide range of environments in which early massive quiescent galaxies live, ranging from regions in which the closest massive galaxies reside at projected distances larger than $\sim 20$ Mpc from the quiescent galaxy under consideration (top left panel) to regions with many other massive galaxies at relatively small projected distances from the target galaxy (e.g. bottom first and third panels). In a few cases, massive quiescent galaxies appear to reside on clear filamentary structures (e.g. third top panel), towards the center of a void-like structure (e.g. second top panel), or at the knot connecting different filaments (e.g. the bottom right panel - incidentally, this region contains two of the massive quiescent galaxies selected, the second being in the top right corner of the projected box). The star formation rates of the galaxies residing in the projected regions considered ranges from very low levels consistent with zero to $\sim 250\,{\rm M}_{\odot}{\rm yr}^{-1}$, with a non negligible number of massive actively forming stars in each region (see Fig. 12).
Although a direct comparison is difficult, the range of environments visible in Fig. 3 is much broader from that depicted in Fig. 4 of Remus and Kimmig (2025), showing the projected surface brightness of the gas at z=3.4 with superimposed the position of the six most massive quiescent galaxies in the Magneticum Simulation Box3. In this simulation, massive quiescent galaxies at high-z do inhabit different regions of the cosmic web as is also in the case in our model, but none of them is found at the center (or close to the center) of the hottest regions (the most massive halos). As discussed in Section 1, this qualitative difference in the spatial distribution of massive quiescent galaxies is interesting because it suggests that it can be used, in principle, to constraint different physical mechanisms driving galaxy quenching at high-z.
As discussed in previous work (Xie et al., 2024, see also Lagos et al. 2025), in our model, the star formation in massive galaxies is suppressed by AGN driven winds that are primarily driven by mergers. Fig. 4 shows the evolution of the parent halo mass, galaxy stellar mass, star formation rate, and black hole mass for the same massive galaxies at $z\sim 4.5$ considered in Fig. 3 (green and magenta lines are used for the lowest and highest percentiles of the o33 distribution, with line thickness and color shade increasing with increasing overdensity value). The figure shows that, already in this relatively small sample of galaxies, there is a large variety of evolutionary histories. The most massive galaxies do not necessarily reside in the most massive halos at the epoch of observation, as already discussed in the previous section, although the stellar mass assembly history parallels that of the parent dark matter halos as expected. The most striking trend as a function of over-density is that galaxies residing in the least over-dense regions tend to sit in halos that have started growing earlier than their counter-parts residing in regions that are over-dense at the epoch of the observation. The earlier stellar/halo mass growth translates into an earlier growth of the black hole mass that can lead to an earlier quenching of the star formation. Therefore, in this sub-sample of galaxies, lower over-densities are explained by an accelerated growth of halo/stellar mass before the epoch of the observations. It is important to stress that these galaxies have not been selected randomly, but to emphasize the variety of the environments at these early cosmic epochs. We will show later that the trend just discussed becomes insignificant when considering large samples of galaxies and large-scale overdensities as a proxy for the environment. However, it remains significant when considering local densities.
A more quantitative characterization of the environment of the quiescent galaxies in our sample is given in Fig. 5 that shows the correlation between two different estimates of the galaxy environment and the parent halo mass (left column), the intrinsic galaxy stellar mass (middle column), and the same quantity convolved with an uncertainty of $0.25$ dex (right column). The top panels correspond to the local density estimated considering the three closest neighbors for each model galaxy (n33), while the bottom panels correspond to an over-density estimated over a $3$ Mpc scale (o33). Black and cyan lines show the distributions for quiescent and star forming galaxies, respectively. The numbers in the top-left and bottom-right of each panel correspond to the Spearman’s correlation coefficient and relative significance (a small value indicates a significant correlation) for the two galaxy samples. The top-left panel shows that there is a good although broad correlation between n33 and the parent halo mass, both for star forming and quiescent galaxies. A weak correlation is also visible between n33 and the intrinsic stellar mass (top-middle panel): the most massive galaxies tend to have larger local densities, but the scatter is very large and large values of n33 can be found for a quite broad range of galaxy stellar masses. This already weak correlation is almost entirely washed out when including an (optimistic) estimate of the uncertainty on the galaxy stellar mass (top-right panel). The correlations with o33 are overall weaker and an uncertainty on galaxy stellar mass can even invert the correlation with galaxy stellar mass (bottom-right panel). The trends just discussed are very similar for quiescent and star forming galaxies, but for a small shift towards lower halo and stellar masses for the star forming galaxies. This shift appears to be more evident when considering local densities than larger scale over-densities. Results are qualitatively similar when considering local densities based on a larger number of neighbors or over-densities computed on larger scales.
5 Do physical properties of high-z quiescent galaxies depend on the environment?
Figure 6: Formation time (top panel) and [O/Fe] (bottom panel) as a function of galaxy stellar mass. The former is defined as the age of the Universe at which half of the stars in the galaxies are formed. Color coding is as a function of redshift, as indicated in the legend in the bottom panel. Open and filled symbols with error bars in the top panel correspond to observational measurements by Carnall et al. (2024, 5 galaxies between $z\sim 3.2$ and $z\sim 4.6$) and Nanayakkara et al. (2025, 19 galaxies with a median redshift of $\sim 3.55$), respectively. Contours in the bottom panel show the regions enclosing (from thicker to thinner) 30, 60, and 90 per cent of the model sample.
Having characterized the environments in which massive quiescent galaxies live, we can then ask to what extent their physical properties depend on their local or global environment. In this section we focus on two different quantities for which observational measurements are rapidly being collected and/or will rapidly increase in the future: star formation histories and the stellar metallicities.
Figure 7: The star formation histories of quiescent galaxies in the bottom 25th (cyan) and highest 75th (magenta) percentiles of the distribution of n33, considering matched distributions of parent halo mass. Thicker lines show the median star formation histories for each sub-sample. The numbers of galaxies in each sub-sample at each snapshot is given in the top left corner. A maximum of 50 individual star formation histories in each sub-sample is shown to avoid overcrowding the panels.
Galaxies in our sample of model quiescent systems are old, metal-rich, and $\alpha$-enhanced. Fig. 6 shows the distributions of the formation times, defined as the age of the Universe when half of the stellar mass has been formed [^7] For our model galaxies, this is computed by linearly interpolating the predicted stellar mass at the outputs available., and the [O/Fe] abundances as a function of galaxy stellar mass. The color-coding indicates the redshift (as in Fig. 2). Open symbols with error bars in the top panel correspond to the observational estimates by Carnall et al. (2024) for 5 massive quiescent galaxies between $z\sim 3.2$ and $z\sim 4.6$. Filled symbols correspond to measurements by Nanayakkara et al. (2025), based on a sample of 19 quiescent galaxies with a median redshift of $\sim 3.55$. For all these galaxies, JWST/NIRSpec spectroscopy is available. The two pairs of symbols connected by dashed lines correspond to the galaxies: ZF-UDS-6496 at $z\sim 4$, with an estimated formation redshift of $\sim 5.6$ in Carnall et al. (2024) and $\sim 5.9$ in Nanayakkara et al. (2025), and ZF-UDS-7329 at $z\sim 3.2$ with an estimated formation redshift of $\sim 11.2$ in Carnall et al. (2024) $\sim 8.8$ in Nanayakkara et al. (2025). The range of estimated formation times for the same systems clearly shows the large systematic uncertainties, potentially more relevant than the reported statistical errors, associated with these measurements. One obvious source of systematic uncertain is a different (e.g. more top-heavy) stellar initial mass function - this is something that we plan to address in future work.
[…]
Fig. 7 shows the average star formation histories of the galaxies in our sample in the 25th (dark cyan) and 75th (dark magenta) percentiles of the distribution of n33. We have omitted the very few galaxies in the highest redshift snapshots considered because of the low number statistics, and we have matched [^8] Due to the significant correlation between n33 and halo mass (top left panel of Fig. 5, there is still a systematic small difference between the median halo mass in the two distribution of $\sim 0.1$ dex. the distributions of parent halo mass for the galaxies in the two extreme environments considered. The number of galaxies used at each redshift and for each environment bin is shown in the top left corner. The figure shows that there is a clear difference between galaxies residing in the regions corresponding to the highest and lowest values of n33. The trends reflect those described for Fig. 4: star formation tends to start earlier, proceed at larger rates, and declines earlier in regions corresponding to the lowest percentiles of the n33 distribution. No strong difference is observed for the times-scale of quenching, that is in all cases very short (see Fig. 11 in Lagos et al. 2025). We have verified that the differences evident in Fig. 7 are largely unaffected when considering local estimates based on a larger number of neighbors (n55 and n77). The different star formation histories corresponding to galaxies in regions with different (local) densities do not translate into a systematic difference in the ages of the stellar populations (when using the formation time as a proxy) of galaxies residing in these different environments. This is due to the very large scatter of the individual star formation histories as indicated by the thin lines (to avoid overcrowding the panels, we show a maximum of 50 - randomly selected - galaxies in each subsample at each redshift). We have also verified that, when considering a global estimate of the environment, the differences between the median star formation histories corresponding to the two extremes is significantly reduced (see Fig. 13 in Appendix A).
Figure 8: Mass-(stellar)metallicity relation color-coded as a function of local density (n33). The upper panel shows the correlation between stellar metallicity and intrinsic galaxy stellar mass, while in the bottom panel model masses have been perturbed assuming an uncertainty of 0.25 dex on the galaxy mass.
Fig. 8 shows the mass-metallicity relation for all quiescent galaxies in our sample, color coded as a function of log(n33). In the top panel, the intrinsic galaxy stellar mass has been used, while in the bottom panel the galaxy stellar mass includes an uncertainty of $0.25$ dex. The figure shows that there is a very weak trend with local density that is more evident for the galaxies with lowest mass in our sample: stellar metallicity is larger for galaxies in regions of lower local density as a consequence of the more prolonged star formation histories. However, this trend is completely washed out when an (optimistic) uncertainty on the galaxy stellar mass is included. We have verified that no significant trend is visible in case a global density environment is considered. In absolute values, the available observational estimates of stellar metallicity reach larger values than those shown in Fig. 8, but cover a wide range and are affected by large uncertainties: Nanayakkara et al. (2025) report values of stellar metallicity ranging between $\sim-0.55$ and $\sim 0.54$; for three of the galaxies in Carnall et al. (2024), the estimated metallicity is 0.320.350.32-0.35 while for the other two the estimated values are $0.41^{0.06}_{-0.09}$ and $0.96^{0.04}_{-0.09}$. In fact, these measurements for high-redshift quiescent galaxies are very challenging because of the very weak metal absorption features due to young (in absolute terms) ages and of the highly $\alpha$-enhanced stellar populations. We refrain to add observational estimates to Fig. 8 because of the very large estimated uncertainties, but note that model predictions also cover a wide range of metallicities.
6 What are the descendants of early quiescent galaxies?
One last question we can ask, taking advantage of our model results, is related to what are the descendants of high-z quiescent galaxies and if their future fate depends on the environments in which they reside. As noted above, about 88 per cent of the galaxies selected at high-z are centrals. Tracing their descendants down to $z=0$, we find that about 77 per cent are still centrals while about 11 per cent are orphans. Fig. 9 shows the distribution of parent halo mass (top panel) and stellar mass (middle panel) both for the galaxies selected at high redshifts (solid lines) and their descendants at $z=0$ (dashed histograms). The purple and cyan lines correspond to the galaxies that are in the extremes (15th and 85th) percentiles of the distribution of o33 at high-z. The bottom panel shows the correlation between the parent halo mass at $z=0$ and the over-density o33 at the time of observation. The figure shows clearly that the distributions of halo masses and stellar masses of the $z=0$ descendants of high-z quiescent galaxies are very broad, broader than the distribution found at the redshift of observation. This can be understood as a natural consequence of the stochasticity of merger events. The effect, that is stronger the higher the redshift, has been noted in earlier work and recognized as a generic prediction of hierarchical growth (e.g. De Lucia and Blaizot, 2007; Trenti et al., 2008; Angulo et al., 2012). Fig. 9 also shows that the large-scale overdensity has a significant impact on the fate of the high-z quiescent galaxies: galaxies that are sitting in the largest overdensities have higher probabilities to end up in a massive cluster today. There is a clear and rather strong correlation between the over-density at the time of the observation and the halo mass at $z=0$, which makes the latter a good predictor of the environment in which the descendants of high-z passive galaxies live at present. A clear but less significant difference appears when considering a local estimate of the environment. The latter is a more powerful indicator of the halo mass at the redshift of observations, but a less powerful indicator of the $z=0$ halo mass (see Fig. 14 in Appendix A). Similar results have been found in the recent work by Remus and Kimmig (2025). Their Fig. 15 shows a broad correlation between more dense environments at $z=3.4$ and larger $z=0$ halo masses. However, 75 per cent of their high-z quiescent galaxies lie in underdense environment according to their definition. As noted above, this is likely related to how AGN feedback is implemented in their simulation, and it is worth noting that their particular implementation of stellar and AGN feedback does not lead to a good agreement with the observed evolution of the galaxy stellar mass function at lower redshifts (Hirschmann et al., 2014; Steinborn et al., 2015).
Figure 9: Solid histograms show the distribution of parent halo mass (top panel) and stellar mass (middle panel) for all quiescent model galaxies. Dashed histograms are the corresponding distributions for their descendants at $z=0$. Cyan and magenta are used for galaxies that are in the extremes (15th and 85th percentiles) of the distribution of o3. The bottom panel shows the correlation between the virial mass of the halos in which descendants of galaxies in our sample reside at $z=0$ and the over-density at the time of the observation. Contours in the bottom panel show the regions enclosing (from thicker to thinner) 30, 60, and 90 per cent of the galaxies.
A more detailed characterization of the future star formation history of high-z quiescent galaxies is shown in Fig. 10. For each model galaxy, we have followed the descendant to $z=0$ and identified all snapshots where the sSFR is above the threshold used to identify a galaxy as quiescent ($0.3\times t^{-1}_{\rm Hubble}$ - see Section 2). We have merged together adjacent snapshots when counting these “rejuvenation” episodes. The top panels of Fig. 10 show the distributions of: the number of rejuvenation episodes, the SFR during each episode, and the associated time-scale. The bottom panels show the distributions of: the total number of mergers in the last snapshot just before the rejuvenation episodes, the number of mergers with a mass ratio of at least 1 to 10, and the associated increase of stellar mass (excluding the stellar mass that is accreted through mergers). Magenta and cyan lines are for the 15th and 85th percentiles of the o33 distributions, as in previous figures.
Figure 10: Top panels: distributions of the number of rejuvenation episodes, of the SFR during each episode, and of the associated time-scale. Bottom panels: distributions of the total number of mergers in the last snapshot just before the rejuvenation episodes, of the number of mergers with a mass ratio of at least 1 to 10, and of the associated fractional increase of stellar mass (excluding the stellar mass that is accreted through mergers). Magenta and cyan are for the 15th and 85th percentiles of the o33 distributions.
In the framework of our model, rejuvenation is not a common event in the future lifetime of a massive quiescent galaxy at high-z. About 36 per cent of the quiescent galaxies in the selected sample remain passive down to $z=0$, experiencing no rejuvenation events. For the rest of the galaxies in the sample, the mean number of rejuvenation episodes is $\sim 1.6$ (with a median number of 1). Only very few galaxies experience a number of rejuvenation episodes larger than 4 (15 galaxies in total out of 4788). The overall fraction of galaxies experiencing rejuvenation in our model is larger than quoted in Szpila et al. (2025) but very close to those estimated in Remus and Kimmig (2025). However, we note that these numbers are very sensitive to the exact definitions adopted. The galaxies with more frequent episodes of rejuvenation events tend to reside in the less over-dense regions at the time of observations, but the number statistics are very low.
About 34 per cent of the rejuvenation episodes are associated with a star formation rate lower than $10\,{\rm M}_{\odot}\,{\rm yr}^{-1}$, but $\sim 5$ per cent of the galaxies in our sample have rejuvenation episodes with associated rates larger than $100\,{\rm M}_{\odot}\,{\rm yr}^{-1}$. A larger number of these episodes are associated with galaxies that reside in the over-dense regions at the epoch of observations. The vast majority of the rejuvenation episodes correspond to events occurring over one snapshot only of the simulation: the interval between subsequent snapshots varies between $\sim 110$ and $\sim 270$ Myr at $z\sim 5$ and $z=0$ respectively. Only a few episodes last much longer, i.e. the galaxy accretes significant amount of gas to sustain star formation over several snapshots. Rejuvenation episodes are typically associated with recent mergers. A few of the galaxies in our sample experience relatively large number of mergers, but many of these have very low mass-ratio (see bottom left and bottom middle panel). For about 15 per cent of the galaxies in the sample, the increase in mass associated with the rejuvenation episodes (subtracting the mass accreted through mergers) is non negligible (larger than $10^{10}\,{\rm M}_{\odot}$). The fractional mass increase with respect to the present stellar mass tends to be larger for galaxies in underdense environments (bottom-right panel), due to the large difference between the present day stellar mass of galaxies residing in overdense and underdense environments (see middle panel of Fig. 9).
7 Summary and conclusions
In this work, we have analyzed the environment of massive quiescent galaxies at $3\lesssim z\lesssim 5$ using our GAlaxy Evolution and Assembly (GAEA) theoretical model of galaxy formation. We selected galaxies with stellar mass larger than $10^{10.8}\,{\rm M}_{\sun}$ and specific star formation rate below $0.3\times t^{-1}_{\rm Hubble}$, resulting in a sample of about $5,000$ galaxies within a simulated volume of $\sim 685$ Mpc comoving on a side. In previous work, we have shown that the latest rendition of our model reproduces a wide range of observational results, including the observed quiescent fractions and number densities up to $z\sim 3-4$, providing an ideal tool for the analysis presented in this paper. We show that the formation times derived for model galaxies at $z\gtrsim 3$ cover well the range inferred from recent observational work (Carnall et al., 2024; Nanayakkara et al., 2025). Galaxies with very short formation time-scales and early formation epochs are present in our model, though they are rare. Systems like ZF-UDS-7329 are therefore not impossible in the framework of our model, though their true abundance in the Universe remains uncertain.
Model high-z quiescent galaxies are $\alpha$-enhanced and exhibit a wide range of stellar metallicity broadly consistent with observational trends, though they do not reach the highest metallicities recently inferred. These measurements remain challenging due to the weakness of diagnostic spectral lines and the lack of well tested $\alpha$-enhanced stellar population models for young ($\lesssim 1$ Gyr) ages. More accurate observational constraints will be essential to use these chemical signatures to test and refine current galaxy formation model.
Our main conclusions can be summarized as follows:
(i) Environmental diversity: massive quiescent high-z galaxies inhabit a wide range of environments, spanning from underdense regions resembling voids, to filaments, walls, and dense knots at the intersections of filaments. Their environment, whether characterized by halo mass, local neighbors counts, or large-scale overdensity, are far from uniform.
(ii) Environmental trends and halo history: quiescent galaxies that live in under-dense regions at the epochs of the observation tend to reside in halos that collapsed earlier and experienced rapid early growth. This trend is most evident when considering a local estimate for the environment, and becomes less clear when using large-scale overdensity. In addition, these correlations can be diluted or even inverted if halo mass is not controlled for, making their observational detection challenging given current uncertainties and the intrinsic large scatter in star formation histories.
(iii) Descendant diversity: the descendants of high-z massive quiescent galaxies span a broad range of halo and stellar masses by $z=0$, due to the stochastic nature of mergers. Some galaxies evolve little and remain in group-size halos, while others grow significantly and end up in cluster-size halos by $z=0$. The large-scale overdensity measured at the epoch of the observations serves as a useful, though not perfect, predictor of the descendant halo mass at $z=0$.
(iv) Rejuvenation: about 36 per cent of the massive quiescent galaxies in our sample do not experience any subsequent rejuvenation, and only a few galaxies undergo multiple (>4>4) rejuvenation events accompanied by strong bursts of star formation. Most rejuvenation episodes are merger-driven and occur more frequently in overdense environments at the epoch of observation.
These findings have important implications for current and future observational programs aimed at using environmental diagnostics to test quenching mechanisms or to identify the progenitors of present-day massive clusters. Both the number densities and environmental properties of massive quiescent galaxies at $z\gtrsim 3$ are expected to exhibit significant cosmic variance. Furthermore, such galaxies do not necessarily reside in the most massive halos of their epoch, nor do they always trace the sites that will evolve into the most massive clusters by $z=0$. This is only partly due to uncertainties in stellar mass estimates and largely reflects the stochastic nature of mass accretion, as discussed in previous work (e.g. De Lucia and Blaizot, 2007; Trenti et al., 2008; Overzier et al., 2009; Angulo et al., 2012, see also Fontanot et al. in preparation).
Acknowledgements.
We acknowledge the use of INAF-OATs computational resources within the framework of the CHIPP project (Taffoni et al., 2020). MH acknowledges funding from the Swiss National Science Foundation (SNSF) via a PRIMA grant PR00P2-193577 ‘From cosmic dawn to high noon: the role of BHs for young galaxies’.
Appendix A Additional figures
Figure 11: As in Fig. 3 but showing in color only galaxies with star formation rates lower than $\sim 5\,{\rm M}_{\odot}\,{\rm yr}^{-1}$.
Figure 12: As in Fig. 3 but showing in color only galaxies with star formation rates larger than $\sim 80\,{\rm M}_{\odot}\,{\rm yr}^{-1}$.
Figs. 11 and 12 show the projected positions of galaxies with star formation rates lower than $\sim 5\,{\rm M}_{\odot}\,{\rm yr}^{-1}$ and larger than $\sim 80\,{\rm M}_{\odot}\,{\rm yr}^{-1}$, respectively. The boxes correspond to those in Fig. 3, with gray symbols showing all galaxies with stellar mass larger than $10^{9}\,{\rm M}_{\odot}$ independently of their star formation rate. A certain number of galaxies forming stars at large rates can be found in all boxes considered. In most cases these tend to avoid the central regions of the boxes but in a few cases, due to projection effects and to the large projected volumes shown, they appear to lie very close to the central regions.
Figure 13: The star formation histories of galaxies in the bottom 25th (cyan) and highest 75th (magenta) percentiles of the distribution of o33. Ticker lines show the median star formation histories. The numbers of galaxies in each sub-sample at each snapshot is given in the top left corner. A maximum of 50 individual star formation histories in each subsample is shown to avoid overcrowding the panels.
Fig. 13 is the equivalent of Fig. 7 but showing the average star formation histories of galaxies in our quiescent sample in the 25th (dark cyan) and 75th (dark magenta) percentiles of the distribution of o33. While there is a clear different (albeit with a very large scatter) between the extremes considering the n33 distribution as discussed in the main body of the paper, this difference is largely removed when considering larger scale overdensities.
Figure 14: As in Fig. 9, but cyan and magenta histograms in the top and middle panel corresponds to galaxies in the extreme of the n33 distribution. The bottom panel shows the correlation between the halo mass in which descendants of quiescent high-z galaxies reside at $z=0$ and n33 at the time of observation.
Fig. 14 is the equivalent of Fig. 9 but for galaxies in the extremes of the n33 distribution. The figure shows that n33 is a less powerful indicator of the $z=0$ halo mass: the scatter shown in the bottom panel is significantly larger than that discussed in Sec. 6 when considering o33. However, the correlation is still present so that galaxies in the extreme of the distribution still have offsetted distributions of virial mass and stellar mass, as shown in the top and middle panels.